Calculate the temperature change that 250. g of water would undergo if 40,000. J were released. (use the specific heat of water: 4.2 J/g oC)

Give your answer in oC.


1
Expert's answer
2021-02-10T01:47:46-0500

Mass= 250g

H= 40000J

C= 4.2J/g°C

∆t= ?

From the formula for quantity of heat

H= mc∆t

∆t= H/mc = 40000/(250x4.2)

∆t= 38.1°C